OpenStudy (anonymous):

In a certain dice game, three fair dice are rolled and the total on all three dice is counted. Let the random variable x be the total count on all three dice. You pay one dollar to roll the dice. If you roll a 16 or greater, you will win $10. If you roll a 6 or lower, you win $5. If you roll anything else, you lose your money. What is the expected value of this game?

OpenStudy (anonymous):

we need to compute the probability of each of these events

OpenStudy (anonymous):

shouldn't be that bad, 16 or higher means a) (5 5 6), (5,6,5) ,(6,5,6) b) 5 6 6 another three c) 6 6 6 only one for a total of 7 out of the possible \(6^3=216\) rolls, probability it \[\frac{7}{216}\] that you win \(\$10\)

OpenStudy (anonymous):

Good job dude.

OpenStudy (anonymous):

six or lower, many more choices @emah, thanks there is lots more to do

OpenStudy (anonymous):

I know, I just also know I couldn't explain it that well.

OpenStudy (anonymous):

here is a picture of the probabilities

OpenStudy (anonymous):

then for the sixes, looks like 1, 1, 1, 1 of these 1, 1, 2 (3 of these) 1, 1, 3 (3 of these) 1,1,4 (3 of these) 1 2 2 (3 of these) 1,2,3 (3 more) 2,2,2 1 of these

OpenStudy (anonymous):

oooh NOW you tell me you have these listed!! then it is real easy

OpenStudy (anonymous):

sorry, i was trying to upload the pic asap lol

OpenStudy (anonymous):

lol^

OpenStudy (anonymous):

i get the second part but not the first one can you guys explain it to me

OpenStudy (anonymous):

then the probability you get a number 6 or less is \[\frac{20}{216}\] probability you get a number 16 or greater is \[\frac{10}{216}\]

OpenStudy (anonymous):

so are we looking for the expected value of rolling a specific number or how much money would be won?

OpenStudy (anonymous):

and you can win \(\$9, \$4\) or \(-\$1\)

OpenStudy (anonymous):

then the expected winnings is the average it is \[9\times P(X=9)+4\times P(X=4)-1\times P(X=-1)\]

OpenStudy (anonymous):

that is \[\frac{9\times 10+4\times 20-1\times 186}{216}\]

OpenStudy (anonymous):

the reason it is 9 and 4, rather than 10 and 5, is presumably you don't win ten dollars plus get your one dollar back

OpenStudy (anonymous):

you can think of it more simply this way imagine you play exactly 216 times and that the probabilities are exact you pay a total of $216 you win back \(10\times 10+5\times 5=100+25=125\) for a loss of \(125-216=-91\)

OpenStudy (anonymous):

averages over each play, you lose \(-\frac{91}{216}\) each time that is your expected value

OpenStudy (anonymous):

@Alessandr09 you here to right? because you had the same question is it clear how to do it?

OpenStudy (anonymous):

oh I see! so would I just leave it as a fraction or would it be better to put it as a decimal?

OpenStudy (anonymous):

that is up to you i guess since it is money, you might want a decimal

OpenStudy (anonymous):

okay, thank you sooooo much!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

thank u i got it @satellite73 you saved my life

OpenStudy (anonymous):

yw, glad to help (and that your life is saved)

OpenStudy (anonymous):

what would the probability distribution look like?

OpenStudy (anonymous):

@aprilmelody, I attached a photo in an earlier reply. It would basically look like that but only for the section with three dice

OpenStudy (anonymous):

ok! could you explain why it goes up to 18?

OpenStudy (anonymous):

or do I stop at 6 rolls?

OpenStudy (anonymous):

You would only stop at 6 rolls if you were only dealing with one die. It goes up to 18 rolls because since there are three dice being rolled, there are a total of 18 outcomes (6+6+6=18).

OpenStudy (anonymous):

So the first two columns, you can just ignore. All you wanna focus on is the third column (3 dice). :]

OpenStudy (anonymous):

Thanks!:DD

OpenStudy (anonymous):

You're welcome :]